Use the elimination method to solve the system of equation.Choose the correct order pair
-x+5y=4 and 4x+3y=16

Answer by Alan3354(69443)

Use the elimination method to solve the system of equation.Choose the correct order pair
-x+5y=4 --> -4x+20y= 16
and 4x+3y=16
-------------------
-4x+20y= 16
4x+3y=16
---------------------------- Add
23y = 32
y = 32/27
---
-x+5y=4
-x + 160/27 = 4
x = 160/27 - 4 = 52/27